The surroundings do 1.44kJ of work compressing a gas in a perfectly insulated cylinder such that no heat can be exchanged. "Calculate the change in internal energy of the system and determine whether the process is endothermic or exothermic"


1
Expert's answer
2021-06-03T05:27:30-0400

ΔE = q + w

q = 0 in a perfectly insulated cylinder such that no heat can be exchanged.

ΔE = 0 + 1.44 = 1.44 kJ

The process is isothermic.